OpenStudy (liizzyliizz):

((2/x^2 )/10)/x^5 (Ok... I've done this several times and I'm not sure which answer I'm getting is correct...)

OpenStudy (anonymous):

\[\frac{2}{x^2}\times \frac{x^5}{10}\] is a start

OpenStudy (anonymous):

then some canceling to get \[\frac{x^3}{5}\]

OpenStudy (liizzyliizz):

I did the beginning right, I must of over analyzed this because i went beyond and above after that... and got something different at first... But thanks :D